MONT 100N – Modeling the Environment Reading/Study Questions on Laudato Si’ September 1,4,8, 2017

Background

Our first reading of the semester in this Montserrat seminar will be the Papal Encyclical Letter Laudato Si’ published by Pope Francis in 2015, which is posted in electronic form on the course homepage. The title comes from the first words of a song attributed to St. Francis of Assisi, “Praise be to you, my Lord,” in which St. Francis expresses his love for everything in the natural world.

Encyclical letters are, traditionally, documents sent by Popes to the bishops and then often relayed to the clergy of the Catholic Church. They deal with some aspect of Church teaching and are used to clarify or amplify some doctrine, or to condemn or promote some issue. Because they are issued through the authority of the Pope to articulate Church doctrine, it is sometimes asserted that they are not to be questioned. Some of you may in fact have heard homilies or participated in discussions about this letter after it was released in 2015.

At the present time, however, encyclical letters are also distributed in print and electronic forms so that they are available directly to all interested people. I think it is now fair to say they (and especially this letter in particular) are often intended in part as ways to provoke discussion among the faithful and others and provide calls to action where some pressing issue needs to be addressed.

To be clear, I do not think we are bound to accept everything in this document just because it comes from a Pope. In fact, I want to be very open and acknowledge that I was not raised in the Catholic Church and I am not particularly religious. However, I do have a tremendous amount of respect for Pope Francis (the first Jesuit Pope!) and I think his point of view in this letter deserves to be heard by everyone.

Whatever beliefs and perspectives you bring to this discussion are welcome, as long as you listen to others, try to understand their points of view, and present your own views in a respectful and civil manner.

We will be reading and discussing Chapters 1, 3, and 5 of the letter in class. I have decided not to ask you to read the other sections because they deal more with the religious justifications for Pope Francis’s ideas. You are free, of course, to read those other sections if you wish.

Some of the following questions are adapted from the study guide from

              https://www3.nd.edu/~pweithma/justice seminar/
                 PopeFrancis/StudyQuestionsonLaudatoSi.pdf

Questions – Chapter 1

1. Much of Chapter 1 is devoted to a fairly detailed discussion of a collection of intertwined environmental issues. Outline what these issues are and how they relate to each other.

2. Do you feel some of the “rapidification” that Francis mentions in §18? Have you ever discussed this with a parent or grandparent?

3. Near the start of Chapter 1, Francis says “Our goal is not to amass information or to satisfy curiosity, but rather to become painfully aware, to dare to turn what is happening to the world into our own personal suffering and thus to discover what each of us can do about it.” (§19) Does what is happening in the world have to become personal suffering for things to change? What kind of attitude toward the world would make that possible? What is the opposite of this?

4. Francis says “access to safe drinkable water is a basic and universal human right, since it is essential to human survival and, as such, is a condition for the exercise of other human rights.” (§30) Is this realistic? What happens if there is not enough to go around? What would it take to ensure that every human being had access to safe drinkable water when water resources don’t necessarily respect national boundaries? (If necessary, look up water resource issues in the Middle East.)

5. One of the most potentially controversial aspects of this letter is the way Francis claims that the environmental problems we are facing and economic inequality are connected. Part of this comes from the fact that “many professionals, opinion makers, communications media and centres of power, being located in affluent urban areas, are far removed from the poor, with little direct contact with their problems.” Would more contact with the lives lived by the poor change anything? What would it take to have more such contact? Is that harder or easier in our current technological society than it was in the past?

6. At the end of Chapter 1, Francis says: “As often occurs in periods of deep crisis which require bold decisions, we are tempted to think that what is happening is not entirely clear. Superficially, apart from a few obvious signs of pollution and deterioration, things do not look that serious, and the planet could continue as it is for some time. Such evasiveness serves as a licence to carrying on with our present lifestyles and models of production and consumption. This is the way human beings contrive to feed their self-destructive vices: trying not to see them, trying not to acknowledge them, delaying the important decisions and pretending that nothing will happen.” Is this a fair characterization of the way things stand in the U.S. at present? What is Francis referring to with the “we are tempted to think that what is happening is not entirely clear?” How can powerful economic interests exaggerate and exploit that uncertainty?

Questions – Chapter 3 (The Human Roots of the Ecological Crisis)

7. How would you characterize Francis’s general view of technology? Of information and communication technology in particular? Has our wisdom in using our technology kept up with our progress in learning how to do things?

8. In §106, what does Francis mean by “an undifferentiated and one-dimensional paradigm” that of “a subject, who using logical and rational procedures progressively approaches and gains control over an external object”? What does he think is wrong about this purely scientific, “technocratic” approach? Is it always wrong to apply that kind of approach to questions about nature? to questions about political and economic life (§109)?

9. Francis argues that the ultimate source of this point of view is what he calls “modern anthropocentrism” (§115). What does he mean by this? How does he see a religious aspect here – a misunderstanding of scripture (at least within Christianity?) What does he think is a healthier way to understand the place of human beings?

10. In §113, Francis says that “people no longer seem to believe in a happy future.” Is he right? Do you think you will be better off economically than your parents are? What do you think the world will be like when you are Prof. Little’s age (61)? If people don’t believe in a happy future, why don’t they? If they don’t believe in a happy future because growth has slowed and middle class wages have stagnated (as is certainly true in the U.S.), would that in fact go against Francis’s point rather than work in favor of it?

11. What does Francis mean by “practical relativism” in §§122 - 123? How is “throwaway culture” and a “disordered desire to consume more that what is really necessary” related to this? Do you agree with this criticism of our consumer culture?

12. Francis sees work as a basic human vocation, or calling. But this is deeper than just saying everyone should have a job and make enough money to have a decent lifestyle. What are the dimensions of this human need?

13. What do you think of the Francis’s remarks about business at the end of §129? What would most business people think of them?

Questions – Chapter 5 (Lines of Approach and Action)

14. Francis summarizes some efforts to develop international agreements to address greenhouse gas emissions at the start of this chapter. This was written before the Paris Climate Accords in 2015. (Look these up if you need to.) Do you think those accords would have changed his point of view somewhat? What about more recent developments (e.g. from June 2017)?

15. Francis rejects the “cap and trade” idea in §171. What is his rationale for that rejection? Do you agree with it? Is his general distaste for financial markets leading him down the wrong path here?

16. A longstanding principle of Catholic thought is “subsidiarity” (see §196), according to which problems should be handled at the lowest or most local level competent to handle them. In §175, though, what does Francis seem to say about the level solutions to these problems must come from?

17. In §178, Francis criticizes the short-sightedness of much political decision-making. Isn’t short-sightedness the price we pay for democracy?

18. In §189, the Pope is highly critical of the world’s response to the financial crisis of 2008. Is he right to be so critical? His criticism seems to depend upon a distinction between the real economy and the financial one. What is this distinction? Is it tenable?

19. In §193, Francis seems to suggest that economic growth be contained or reversed in order to achieve “sustainable development.” But economic growth in the developing world has lifted millions into the “global middle class” in recent decades and there is strong sentiment in countries like India that they deserve to experience the benefits of economic development like that that has occurred in North America and Europe. Do more developed countries have the right to say “no” to that? Francis suggests redefining what we mean by “progress” (§194) but is that realistic or desirable?

20. In §195, Francis says “The principle of the maximization of profits, frequently isolated from other considerations, reflects a misunderstanding of the very concept of the economy.” Is the Pope right about this? Consider, in this connection, that the word economy comes from a Greek word meaning “the household” and management of a home. If the earth is, as Francis says, our “common home,” than maybe economics is–literally or at least etymologically– about stewardship of the earth.